# CLIX: Cross-Lingual Explanations of Idiomatic Expressions

### v1

@@ 5.2 A Note on Translation
Although there has been success translating idiomatic expressions in the past (Fadaee et al., 2018; Zaninello and Birch, 2020), translation alone is not sufficient in educational scenarios. In many past approaches, the task assumes that a model will be provided with at least sentence-level context, which helps models perform well on translation-based tasks. However, there is no guarantee that a live user will give a model an entire sentence or more when translating an idiomatic expression. Our task tests the robustness of these models to a lack of context when a user query may only contain an idiomatic expression they don’t understand.
To test this assumption, we ask GPT-3.5 Turbo to provide only a translation of the idiomatic expressions in our test set. When comparing the semantic similarity of the generated translation to the gold explanation, we see that this approach appears to perform better than that of the fine-tuned models used in our experiments in Tab 3, earning scores of 45.80 and 44.21 for Spanish and German respectively. However, we asked our annotators to take a closer look at the outputs to make note of 3 things for each sample; Is the translated expression good/natural? Is the translation done word-for-word? Is this translation an instance of an idiom in your native language?
Using these 3 factors, we analyze how often a sample is above a threshold of 40555We found the scores had incredibly high standard deviations of 20.52 and 19.84 respectively, so we set this threshold at a rough 10% under the mean. while still being judged as a poor or unnatural expression, which was the case for 14.5% and 18.5% of our test set in Spanish and German respectively. While semantic similarity is not meant to represent the actual quality of a sample, this further indicates that the similarity score we obtain cannot be the only metric used for evaluating models on tasks such as these. Furthermore, of the total number of expressions that were found to be good/natural, the proportion that were word-to-word translations were 48.3% and 56.3%, where 71.4% and 93.1% of those subsets were a case where the idiom was also present in the target language. This tells us that translation of just an idiomatic expression is not likely to generalize, as one will often need the same idiom to be present in the target language. These findings also hint at the potential that models trained to capture multilingual sentence similarity in their embedding spaces are more strongly capturing word-to-word similarity than meaning similarity in the case of idiomatic expressions.
@@ 6 Human Evaluation
Our annotators found that the model performed extremely well in terms of both Fluency and Accuracy, with average scores of 4.70 and 4.78 respectively. To measure inter-annotator agreement, we calculated Krippendorff’s . These scores range from -1 to 1, where 1 is perfect agreement, 0 is random agreement and -1 is inverse agreement. We found substantial agreement for fluency () and moderate agreement for accuracy (). Upon closer inspection, we found two main reasons for disagreement. First, in some cases, an explanation may begin with a literal interpretation and end with a figurative interpretation. In this case, one of our annotators appeared to penalize for the inclusion of the literal interpretation where the other did not. Second, our raters’ native dialects of Spanish are different, resulting in differences in understood meaning or connotation of certain words and phrases.
@@ 6.1 Error Analysis
During human evaluation, our raters pointed out common error types affecting both fluency and accuracy. Issues with fluency come from the model’s tendency to repeat words that may be key to the meaning of the idiomatic expression or make grammatical mistakes. Problems with accuracy arise in cases where the model is either too literal or misidentifies the specific type of entity or domain that an expression refers to. Another observation made by our annotators was that the model might introduce information not necessarily related to the meaning of the idiomatic expression. Rather, it might indicate the origin of the idiomatic expression, or an indicated stance of the speaker. We do see this as a positive addition to the explanation in educational contexts, but there is inconsistency and we neither ask for nor measure this phenomenon in our experiments. In the context of our educational goals, these findings present a few key issues that we would not wish to see arise in classroom settings, for example. Despite our high fluency and accuracy scores reported in Sec. 6, there is clearly still room to grow before these models are given real responsibility in educational domains.
@@ 7 Conclusion and Future Work
The primary limitation of our paper is the small size of our dataset. The fact that we only have access to 628 idioms (resulting in 1,256 examples total, one for each target language) limits our ability to learn a good generation model. To combat this issue, we wish to augment our dataset with additional idioms, using resources such as the Oxford Dictionary of Idioms, which includes a total of 1,506 idioms. Additionally, are limited in the number of languages we support. We have only one source language of idiomatic expressions, English. There is a reasonable chance that the models we have experimented with might perform worse in experiments conducted with languages that have smaller amounts of available data. Furthermore, we only attempt cross-lingual explanation into two languages, Spanish and German, which are similarly well supported with data. Finally, the metrics we use in this paper are simply not able to fully capture the phenomenon we wish to measure. We hope to motivate and/or participate in the development of more informative metrics that best suit a task such as ours.
@@ Appendix A Hyperparameters and Other Tables
For T5 and mT5, we use a learning rate of 1e-5, the AdamW optimizer, a batch size of 32, and run for 1000 epochs with early stopping. For GPT and Llama, we also explore two different configurations to constrain the length of our explanation to a maximum of one or two sentences. Additionally, for GPT 3.5 Turbo we set the temperature to 0.5 for all scenarios, and Llama was left with its default temperature of 0.8. All hyperparameters are chosen using the development set.
@@ Appendix B Human Evaluation Guide
We provide the rubrics given to our annotators for rating both fluency and accuracy in Tabs. 9 and 10.
@@ Appendix C Themes
We include the full list of themes and their frequencies within our dataset (in descending order) in Tabs. 11 and 12.

### v2

@@ 5.2 A Note on Translation
Although there has been success translating idiomatic expressions in the past (Fadaee et al., 2018; Zaninello and Birch, 2020), translation alone is not sufficient in educational scenarios. In many past approaches, the task assumes that a model will be provided with at least sentence-level context, which helps models perform well on translation-based tasks. However, there is no guarantee that a live user will give a model an entire sentence or more when seeking to understand an idiomatic expression. Our task tests the robustness of these models to a lack of context. In Appendix F, we outline an experiment where data from human evaluators indicates that 42% of Spanish and 48.5% of German translations remain unnatural, while good translations will be word-for-word in roughly 50% of cases and thus cannot be used as a viable method for this task unless for an idiomatic expression in there is an equivalent expression in .
5.3 Data Contamination
There is always potential for data contamination in LLM experiments. While we cannot make claims about what data the LLMs may have seen in training, we attempt to correlate the explanation generation performance with the number of times an idiom appears online and the likelihood of generating a particular expression from in context examples using an LLM. We find there is no correlation (Spearman’s between -0.03 and -0.3) between either proxy measure and model performance. We include a comprehensive report in Appendix G.
@@ 6 Human Evaluation
Our annotators found that the model performed extremely well in terms of both Fluency and Accuracy, with average scores of 4.70 and 4.78 respectively. To measure inter-annotator agreement, we calculate Krippendorff’s . These scores range from -1 to 1, where 1 is perfect agreement, 0 is random agreement and -1 is inverse agreement. We found for fluency and for accuracy. We found two main reasons for disagreement. First, in some cases, an explanation may begin with a literal interpretation and end with a figurative interpretation. In this case, one of our annotators appeared to penalize for the inclusion of the literal interpretation where the other did not. Second, our raters’ native dialects of Spanish are different, resulting in differences in understood meaning or connotation of certain words and phrases.
@@ 6.1 Error Analysis
During human evaluation, our raters pointed out common errors affecting both fluency and accuracy. Issues with fluency come from the model’s tendency to repeat words that may be key to the meaning of the idiomatic expression or make grammatical mistakes. Problems with accuracy arise in cases where the model is either too literal or misidentifies the specific type of entity or domain that an expression refers to. Our annotators also observed that the model might introduce information not necessarily related to the meaning of the expression. Rather, it might indicate the origin of the idiomatic expression, or an indicated stance of the speaker. We see this as a positive addition to the explanation in educational contexts, but there is inconsistency and we neither ask for nor measure this phenomenon in our experiments. In the context of our educational goals, these findings present a few key issues that we would not wish to see arise in classroom settings. Despite our high fluency and accuracy scores reported in Sec. 6, there is clearly still room to grow before these models are given real responsibility in educational domains.
@@ 7 Conclusion and Future Work
Given that we cannot make Oxford-ME public, the primary limitation of our paper is the small size of our publishable dataset. Additionally, we are limited in the number of languages we support. We have only one source language of idiomatic expressions, English. There is a reasonable chance that the models we have experimented with might perform worse in experiments conducted with languages that have smaller amounts of available data. Furthermore, we only attempt cross-lingual explanation into two languages, Spanish and German, which are similarly well supported with data. Finally, the metrics we use in this paper are simply not able to fully capture the phenomenon we wish to measure. We hope to motivate and/or participate in the development of more informative metrics that best suit a task such as ours.
@@ Appendix A Hyperparameters and Other Tables
For T5 and mT5, we use a learning rate of 1e-5, the AdamW optimizer, a batch size of 32, and run for 1000 epochs with early stopping. For GPT and Llama, we also explore two different configurations to constrain the length of our explanation to a maximum of one or two sentences. Additionally, for GPT 3.5 Turbo we set the temperature to 0.5 for all scenarios, and Llama was left with its default temperature. All hyperparameters were chosen using the development set.
| Model | Overall | Spanish | German | |||
| BLEU | ROUGE | BLEU | ROUGE | BLEU | ROUGE | |
| mT5 Direct | 0.37 | 9.94 | 0.34 | 10.59 | 0.47 | 9.34 |
| mT5 Pipeline | 0.09 | 4.90 | 0.18 | 9.71 | 0.29 | 7.80 |
| T5 Pipeline | 0.89 | 11.95 | 1.03 | 13.09 | 0.74 | 10.73 |
| Llama Direct Zero-Shot | 1.00 | 12.31 | 1.12 | 14.19 | 0.87 | 10.48 |
| Llama Pipeline Zero-Shot | 1.40 | 14.69 | 1.67 | 16.54 | 1.08 | 12.88 |
| GPT Direct Zero-Shot | 3.58 | 22.1 | 3.66 | 24.39 | 3.50 | 19.87 |
| GPT Pipeline Zero-Shot | 4.59 | 26.41 | 5.56 | 28.21 | 3.57 | 24.63 |
| GPT Direct 5-Shot | 4.04 | 23.23 | 4.22 | 25.09 | 3.86 | 21.34 |
| GPT Pipeline 5-Shot | 4.19 | 24.76 | 4.88 | 26.38 | 3.47 | 23.13 |
@@ Appendix B Human Evaluation Guide
We provide the rubrics given to our annotators for rating both fluency and accuracy in Tabs. 10 and 11. The fluency rubric is inspired in-part by ACTFL (2024).
@@ Appendix C Themes
We include the full list of themes from Rafatbakhsh and Ahmadi (2019) and their frequencies within our dataset (in descending order) in Tabs. 12 and 13.
@@ Appendix D Example Prompts
| Fill-in-the-Blank | • System: Fill in the appropriate idiomatic expression that completes the sentence below. • User: {context} Response: [blank] can be replaced with: |
| Oxford-ME Context Generation | • Use the idiom {idiom} meaning {explanation} in a sentence. Respond with nothing but the sentence. |
| Statistic | EPIE-ME | Oxford-ME | |
| Spanish | German | Spanish | |
| Mean | 0.159 | 0.162 | 0.069 |
| Min | 0 | 0.092 | 0 |
| Max | 0.781 | 0.857 | 0.8 |
| Stdev | 0.162 | 0.216 | 0.126 |
| Q1 | 0 | 0.092 | 0 |
| Q2 | 0.131 | 0.269 | 0 |
| Q3 | 0.249 | 0.440 | 0.095 |
| Q4 | 0.781 | 0.857 | 0.8 |
Appendix E Noise in Translations
Where EPIE-ME needed a small amount of correction, Oxford-ME required significantly less modification.555For Oxford-ME, we verify only the Spanish translations for this analysis. We believe this is primarily due to the difference in English explanation length between the two sources of data. Examining the quartiles of normalized edit distance, we see that 75% of the data required roughly less than a 25% modification to the longer string for Spanish explanations in EPIE-ME, less than a 45% modification for German in EPIE-ME, and less than a 10% modification for Spanish explanations in Oxford-ME. Overall, quite low. The high standard deviations can be explained by the lack of samples near the mean, rather a large portion of samples are in the lower quartiles. These measurements can be found in Tab. 15. Graphs of the same information can be found in Fig. 1.
Appendix F Translation as a Method
To test our assumption that translation is not a wholesale solution, we ask GPT-3.5 Turbo to provide only a translation of the idiomatic expressions in our test set. When comparing the semantic similarity of the generated translation to the gold explanation, we see that this approach appears to perform better than that of the fine-tuned models used in our experiments in Tab 3, earning scores of 45.80 and 44.21 for Spanish and German respectively. However, we asked our annotators to take a closer look at the outputs to make note of 3 things for each sample; Is the translated expression good/natural? Is the translation done word-for-word? Is this translation an instance of an idiom in your native language?
Using these 3 factors, we analyze how often a sample is above a threshold of 40666We found the scores had incredibly high standard deviations of 20.52 and 19.84 respectively, so we set this threshold at a rough 10% under the mean. while still being judged as a poor/unnatural expression, which was the case for 14.5% and 18.5% of our test set in Spanish and German respectively. While semantic similarity is not meant to represent the actual quality of a sample, this further indicates that the similarity score we obtain cannot be the only metric used for evaluating models on this task. Of the total number of expressions that were good/natural, the proportion that were word-to-word translations were 48.3% and 56.3%, where 71.4% and 93.1% of those subsets were a case where the idiom was also present in the target language. This tells us that translation of just an idiomatic expression is not likely to generalize, as one will often need the same idiom to be present in the target language. These findings also hint at the potential that models trained to capture multilingual sentence similarity in their embedding spaces are more strongly capturing word-to-word similarity than meaning similarity in the case of idiomatic expressions.
| Correlation with Similarity Scores | ||
| Model / Proxy | -b | |
| GPT vs SR | -0.181 | -0.123 |
| Llama vs SR | -0.293 | -0.205 |
| Llama vs GP | -0.030 | -0.021 |
| Correlation with Human Evaluation | ||
| Metric / Proxy | -b | |
| Fluency vs SR | -0.141 | |
| Accuracy vs SR | -0.067 | |
| Fluency vs GP | -0.159 | |
| Accuracy vs GP | 0.072 | |
Appendix G Data Contamination
Some may be skeptical of our experiments and the potential for data contamination in our LLMs. Specifically, that they may have seen the idiom-explanation pairs within our dataset. While we do not have access to the training data of either GPT or Llama, we will assume that they have seen almost the entire internet for the purposes of determining the level of contamination. Under this assumption, if the model is performing well on certain idiomatic expressions, we would expect those to be the most used across the internet. For both GPT and Llama, we attempt to correlate both the similarity score and human evaluation results per expression with the number of search results returned by the Google API when using the expression as input – not accounting for morphological shift or interjecting words. For Llama, we additionally extract the probability of generating an expression given 10 in-context examples via fill-in-the-blank prompting.777Prompts found in Tab. 14 Our similarity score results are correlated with the search results and generation probabilities using Spearman’s rank correlation coefficient () and Kendall’s rank correlation coefficient, variant b (-b). Because of the many ties in our human evaluation results, we correlate only with -b. In every single case, we see a score very close to 0, indicating no association whatsoever (Tab. 16). We further examine our results by quartile of search result hits. There is no clear upward trend as search results increase. In fact, the Llama results show an inverse relationship, the Fluency results decrease until the final quartile, and the GPT and Accuracy results show a bell curve peaking in the second quartile (Tab. 17).
| Metric | Q1 | Q2 | Q3 | Q4 |
| GPT Similarity | 71.76 | 72.60 | 69.71 | 67.62 |
| Llama Similarity | 63.40 | 61.99 | 60.85 | 55.81 |
| Fluency | 4.85 | 4.70 | 4.58 | 4.67 |
| Accuracy | 4.78 | 4.86 | 4.80 | 4.69 |